Changes between Version 73 and Version 74 of Status


Ignore:
Timestamp:
Jun 21, 2012 3:17:26 PM (6 years ago)
Author:
emca
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • Status

    v73 v74  
    55Please contact Dr Gary Llewellyn (Science / operations coordinator) if you have any questions or concerns: gaew@nerc.ac.uk (01452 859945 or 07919 697851)
    66
    7 == ARSF Sitrep 07.06.11 ==
     7== ARSF Sitrep 07.06.12 ==
    88''During the rest of June the ARSF intend to collect data from the North of London (RG12-10), a location near Madrid, Spain (CR12-01), Harnhill & Diddington (GB12-06)  as well as several UK based projects that we were hopeing to support in May (GB09-04, GB12-10, BGS08-01 & BGS08-03/04/12-01). Of these RG12-10 holds the highest priority, though as always we are dependent on the weather!
    99''
    1010
    11 == ARSF Sitrep 23.05.11 ==
     11== ARSF Sitrep 23.05.12 ==
    1212''Late on Sunday (20.05.12), Carl (Chief Pilot and Acting Head of the Facility) had to urgently return to the UK on a commercial flight. This related to the health of a member of his family, we wish him the very best of luck. However, James Johnson (our instrument operator) is also a commercial pilot which allowed our transit with the Dornier to continue.'' 
    1313
     
    1919 
    2020
    21 == ARSF Sitrep 21.06.11 ==
     21== ARSF Sitrep 21.05.12 ==
    2222''After two days (Friday and Saturday) of overcast conditions and rain Sunday was sunny and clear. The remaining flight lines over Milos were completed in the morning and the remaining Santorini site was completed in the afternoon. Indications are that all the data (VNIR, SWIR, LIDAR and digital photography) were good.''
    2323
    2424''On Monday we will set off early to Athens where all copies of the data will be transfered to the Greek Military. This transfer is likely to take at least one day.''
    2525
    26 == ARSF Sitrep 18.06.11 ==
     26== ARSF Sitrep 18.05.12 ==
    2727''Yesterday (17.05.12), 11 survey lines (including a cross-line) were flown over the island of Milos in support of EM10-02. This constituted a third of the survey. It was suspended due to a build up of cloud and the last line will be repeated when we resume the survey later in the campaign. 165.8GB of data were collected. These included VNIR, SWIR, LiDAR and digital photography. We arrived on site at 10:27 hrs (local time) and the illumination is such that for most of the survey the Eagle data (VNIR) have a bandwidth of 1.25 nm (the best we have ever achieved). ''
    2828